# EyeDentify: A Dataset for Pupil Diameter Estimation based on Webcam Images

### v1

@@ Abstract
In this work, we introduce EyeDentify, a dataset specifically designed for pupil diameter estimation based on webcam images. EyeDentify addresses the lack of available datasets for pupil diameter estimation, a crucial domain for understanding physiological and psychological states traditionally dominated by highly specialized sensor systems such as Tobii. Unlike these advanced sensor systems and associated costs, webcam images are more commonly found in practice. Yet, deep learning models that can estimate pupil diameters using standard webcam data are scarce. By providing a dataset of cropped eye images alongside corresponding pupil diameter information, EyeDentify enables the development and refinement of models designed specifically for less-equipped environments, democratizing pupil diameter estimation by making it more accessible and broadly applicable, which in turn contributes to multiple domains of understanding human activity and supporting healthcare. Our dataset is available at https://vijulshah.github.io/eyedentify/.
@@ 1 Introduction
The cognitive state of humans is closely linked to features observable through their eyes. Fortunately, the accessibility of eye monitoring in everyday life is rapidly increasing, exemplified by recent advancements such as Apple’s incorporation of camera-based eye tracking features [3, 14]. However, research in this domain primarily targets blink detection [17] and gaze estimation [37, 52], employing various methodologies, including the use of biomarkers [28], infrared spectrum reflected from the eyes [12], or image-based techniques [16]. In comparison, fewer explore pupil diameter estimation [43, 5], which also plays an undeniably crucial role in determining various physiological and psychological states. This oversight highlights a critical gap in the field, underscoring the need for more comprehensive approaches to fully leverage eye monitoring for cognitive state analysis:
@@
Table 1: Comparison of related datasets for eye monitoring. While most datasets have gaze coordinates [17, 54, 9, 22, 53, 18], there is a significant gap in pupil diameter informed [41, 5] datasets.
Dataset Participants Amount of data [frame] Public Gaze Coordinates Pupil Diameter
MAEB [17] 20 1,440
MPIIFaceGaze [54] 15 213,659
Dembinsky et al. [9] 19 648,000
Gaze360 [22] 238 172,000
ETH-XGaze [53] 110 1,083,492
VideoGazeSpeech [18] unknown 35,231
Ricciuti et al. [41] 17 20,400
Caya et al. [5] 16 unknown
EyeDentify (ours) 51 212,073
The third challenge is the prediction of pupil diameter itself. Previous studies [22, 53, 4] have highlighted that estimating gaze coordinates with a camera involves analyzing images of approximately $60\times 36$ pixels [54]. The scale of our images will be smaller for pupil diameter estimation, necessitating analysis at an even finer resolution. This makes accurately predicting pupil diameters more complex than gaze estimation and presents a challenging task.
@@
    Creation of the Pupil Diameter Dataset EyeDentify: We use the Tobii eye-tracker to collect accurate ground truth pupil diameter measurements. Concurrently, we captured facial video recordings with a built-in webcam. In total, data from 51 participants were gathered, comprising 212,073 images (i.e., blinking eyes removed from the initial 226,912 images).
@@
    Data-Collecting Application ChameleonView: Our own implemented web application called ChameleonView captures various pupil diameters. It displays various screen colors during webcam recordings to ensure a diversity of pupil diameters. Moreover, it captures the timestamp and records videos while participants look at the display.
@@
    Initial Pupil Diameter Estimations: We evaluated several models to estimate pupil diameter sizes for both left and right eyes in a 5-fold cross-validation manner. For this, we tested ResNet-18 and ResNet-50, with ResNet-18 emerging as the most accurate, achieving a mean absolute error of 0.1340 $\pm$ 0.0196 and 0.1403 $\pm$ 0.0328 for the left and right eye, respectively.
2 Related Work
This section will examine existing datasets for eye monitoring, including gaze estimation and blink detection (Section 2.1). Furthermore, it surveys datasets for pupil diameter estimation (Section 2.2), as well as methods to estimate pupil diameters (Section 2.3). Table 1 compares all datasets appearing in eye monitoring publications, and we will compare our dataset EyeDentify with them.
2.1 Gaze Estimation and Blink Detection Datasets
Various gaze estimation, blink detection, and facial image data have been collected under different conditions, which we want to examine in the following. For instance, MAEB [17] has collected 20 subjects’ images from nine different angles for robust image-based blink detection purposes. The study aims to complement the open-source dataset HUST-LEBW [19] with additional comparisons. Another example is MPIIFaceGaze [54], which has collected a daily life gaze coordinates dataset from 15 subjects. The strength of this dataset is that it has different illumination conditions and a collection of various day conditions, ranging from nine days to three months. Another dataset is presented by Dembinsky et al. [9], which collected a 360-minute dataset of 19 subjects with their corresponding gaze coordinates while speaking and/or listening. Thus, it incorporates different cognitive activities during recording. An alternative data collection setting is presented by Gaze360 [22], which contains 238 subjects with 172,000 frame images. The distinguishing feature of this dataset is that it also contains outdoor data. Moreover, the camera used during data collection captured 360-degree view recordings, thus including the subjects’ environment. Another varying axis during dataset collection is proposed by ETH-XGaze [53], which collected data from 110 subjects with 1,083,492 images under various hardware settings. More specifically, they used 18 high-resolution Canon 250D digital SLR cameras and four Walimex Daylight 250 light boxes. Lastly, VideoGazeSpeech [18] collected 35,231 frames from 29 videos. As a result, they extracted eye gaze data from existing videos; hence, the feature extraction approach varies from that of other studies.
Most mentioned datasets focus on enhancing gaze estimation and blink detection through diverse and complex recording conditions, but none incorporate pupil diameters.
2.2 Pupil Diameter Datasets
To the best of our knowledge, two data datasets have been used in publications for pupil diameter estimation that employ diverse approaches for data collection. For instance, Ricciuti et al. recruited 17 subjects ranging in age from 15 to 74 years and collected 20-second RGB video recordings under varying light intensities using a GoPro Hero 6 camera at a frame rate of 30fps [41]. Similarly, Caya et al. gathered data from 16 subjects aged 20 to 40 years, employing a camera mounted 10 cm from the subject’s face in a controlled setting. Their study quantified a pupil diameter of 20 pixels as equivalent to 5 mm [5].
Despite these efforts, the mentioned works have not made their datasets publicly available. In contrast, our work and its dataset EyeDentify is not only inspired by their effort of collecting the necessary data for pupil diameter estimation but also presents the largest publicly accessible dataset for pupil diameter estimation from RGB images, significantly contributing to the broader field of eye monitoring.
2.3 Pupil Diameter Estimation Approaches
The estimation of pupil diameters has been done in various research studies. For instance, Ni et al. proposed a pupil diameter estimation method called BINOMAP [35]. Their approach uses master and slave cameras as a binocular geometric constraint for the input gaze images. The estimation model they implement is based on Zhang’s algorithm [55], which achieved $0.022\pm 0.017$mm mean absolute error. Similar to our work, Caya et al. uses their own data recording platform for data collection [5]. They captured the participant’s facial images with a fixed camera distance of 10cm from the face. Then, the facial image is sent to Raspberry Pi for the pupil diameter estimation process. This estimation includes a RGB to grayscale conversion, contrast and brightness adjustment, image reshaping, and applying the Tiny-YOLO algorithm [23]. They achieved pupil measurements with a percent difference of 0.58% for the left eye and 0.48% for the right eye.
In summary, previous works have strong limitations in condition dependencies, such as having two cameras or keeping the face at a constant and fixed distance from the camera. Also, the data collection and evaluation must consider various pupil diameter sizes. We aim to address these limitations with the presented work.
3 Methodology
In this section, we describe the hardware (Section 3.1) and application we used for data collection (Section 3.2), the procedure for collecting the data (Section 3.3), as well as how we processed the raw recordings (Section 3.4).
3.1 Data Recording Hardware
The ground truth data for our pupil diameters is captured using the Tobii eye-tracker [1], a remote device capable of recording gaze coordinates and pupil diameters at 90Hz. It separately measures the diameters of the left and right pupils in millimeters with six-decimal precision. For camera-based recordings of the face and eyes, we utilize the built-in webcam of the Microsoft Surface Studio 1, which offers a screen resolution of $1280\times 720$ pixels and a frame rate of 30.0 frames per second.
3.2 Data Collection Application
A distinctive aspect of our data collection methodology is using an application developed explicitly for EyeDentify. We implemented a web application called ChameleonView  [^1] https://chameleon-view.netlify.app/, which we used for our webcam data collection. The web application shows a button in the center of the screen that, when clicked by subjects, triggers a three-second webcam video recording. Additionally, we capture timestamps marking the start and end of each recording. These timestamps are crucial as they ensure synchronization between the webcam video and the Tobii eye-tracker data, as shown in Figure 1. During one data collection period, the subjects were asked to repeat this button-clicking process $50$ times, resulting in $50$ sessions for each participant.
To collect diverse pupil diameter sizes, we changed the computer screen background color [50, 40]. Thus, the 50 recordings had the following background color conditions (with corresponding HEX codes): the first ten recordings had a white display color (#ffffff), and the remaining 40 recordings were: black (#000000), red (#ff0000), blue (#$0$ff), yellow (#ffff00), green (#008000), and gray (#808080), each in turn, five times displayed and a final display of white (#ffffff) for the last ten recordings again.
3.3 Data Collection Procedure
In this study, subjects began by completing a consent form and providing statistical information. Following this initial step, they proceeded to calibrate the Tobii eye-tracker. After calibration, the subjects started using the data collection application outlined in Section 3.2.
Upon completion of the experiment, the data from the Tobii, webcam video recordings, and corresponding timestamps were consolidated.
Figure 1: Data recording flow. Tobii eye-tracker records pupil diameter, and ChameleonView captures facial recordings using a webcam. Facial recordings start when the participant clicks on the button in the center. The start and end timestamp of the recording is collected in order to synchronize the data with an eye-tracker.
3.4 Data Preprocessing
After recording the raw webcam images alongside the measurements of the eye features, we merge them together for the final dataset EyeDentify. This involves two phases: Aligning the recordings (Section 3.4.1) and cropping the eyes (Section 3.4.2).
3.4.1 Aligning the Recordings
As mentioned in Section 3.2, we first synchronized the data collected from the Tobii eye-tracker with the webcam recordings for each participant, as shown in Figure 1.
We utilized a CSV file generated by the eye-tracker, which records various metrics, including the diameters of the right and left pupils, the average diameter of both pupils, and coordinates for both the left and right gaze positions. A separate timestamp CSV file also captured the start and end times for each participant’s webcam recordings. The synchronization involves iterating through these timestamps and selecting all corresponding rows from the Tobii CSV file that match the start and end times specified for each recording segment. Each recording in our data collection app lasts three seconds, corresponding to three distinct timestamps per recording.
Since the video recording frame rate is 30 frames per second, this results in 90 frames per recording from our data collection application ChameleonView. Additionally, the Tobii eye-tracker operates at a frequency of 90Hz, providing 90 data points per second, resulting in 270 data points for each three-second recording.
Figure 2: Data alignment flow of a single recording. To synchronize the 90 frames with the 270 Tobii-captured data points, each metric column is concatenated horizontally across the 90 data points from the three unique timestamps in the Tobii-captured CSV file, followed by computing a row-wise mean.
@@ 3.4.1 Aligning the Recordings
To synchronize the 90 frames with the 270 Tobii-captured data points effectively, we concatenate each metric column separately across the 90 data points from the three unique timestamps (horizontally) in the Tobii-captured CSV file and compute a row-wise mean. This process is repeated for each recording. The first timestamp of the trio is designated as the primary timestamp for that recording, ensuring consistency in data alignment. Subsequently, we extract frames from the video recordings, yielding 90 images - aligned with 90 data points from the Tobii-eye tracker. This procedure is meticulously repeated for each participant to ensure uniformity and precision in the dataset preparation process. The entire data alignment flow is illustrated in Figure 2.
Figure 3: Pipeline of our data preprocessing. For face detection and landmark localization, we used Mediapipe to extract the respective cropped eye images (32x16), left and right, separately. Next, we applied blink detection on the cropped eyes using the Eye Aspect Ratio (EAR) and a pre-trained vision transformer for blink detection. Cropped eye images are then saved based on the EAR threshold and model confidence score.
3.4.2 Cropping the Eyes
Once the frames are aligned with the Tobii eye-tracker data, the subsequent step involves cropping the eyes from each extracted frame. We used Mediapipe [30] for its robust face and eye detection capabilities. This tool not only identifies faces within the frames but also precisely locates facial landmarks, including the eyes, nose, iris, eyebrows, and mouth. For our purposes, we specifically utilize Mediapipe to crop the left and right eyes.
We utilize Mediapipe’s face landmark detection bundle [2], which includes a series of models designed for face detection and the precise mapping of facial landmarks. The face mesh model identifies 478 distinct facial landmarks and accurately determines the coordinates for the left and right eyes and irises. Given the variability in image intensity and participants’ distance from the webcam, the regions defined by these coordinates may vary in size, leading to inconsistencies in the dimensions of the cropped regions for each eye and iris. To standardize our data, we crop a region around each eye by centering on the detected eye region and fixing the dimensions to 32 pixels in width and 16 in height. This method preserves the natural shape and scale of the eye, avoiding distortions that could result from resizing.
To improve the quality of the set of cropped images, we excluded frames where participants were blinking. For this reason, we first calculated the Eye Aspect Ratio (EAR), a concept adapted from Soukupova et al. [45] and tailored to suit the coordinates provided by Mediapipe’s landmark detection outputs. This adaptation ensures the precise detection of blinks using the specific facial landmarks identified by Mediapipe and is described by

### v2

# PupilSense: A Novel Application for Webcam-Based Pupil Diameter Estimation

@@ Abstract
Measuring pupil diameter is vital for gaining insights into physiological and psychological states — traditionally captured by expensive, specialized equipment like Tobii eye-trackers and Pupillabs glasses. This paper presents a novel application that enables pupil diameter estimation using standard webcams, making the process accessible in everyday environments without specialized equipment. Our app estimates pupil diameters from videos and offers detailed analysis, including class activation maps, graphs of predicted left and right pupil diameters, and eye aspect ratios during blinks. This tool expands the accessibility of pupil diameter measurement, particularly in everyday settings, benefiting fields like human behavior research and healthcare. Additionally, we present a new open source dataset for pupil diameter estimation using webcam images containing cropped eye images and corresponding pupil diameter measurements.
@@ 1 Introduction
The cognitive state of humans is closely linked to features observable through their eyes. Fortunately, the accessibility of eye monitoring in everyday life is rapidly increasing, exemplified by recent advancements such as Apple’s incorporation of camera-based eye tracking features [4, 18]. However, research in this domain primarily targets blink detection [21] and gaze estimation [43, 59], employing various methodologies, including the use of biomarkers [36], infrared spectrum reflected from the eyes [14], or image-based techniques [20]. In comparison, fewer explore pupil diameter estimation [51, 9], which also plays an undeniably crucial role in determining various physiological and psychological states. This oversight highlights a critical gap in the field, underscoring the need for more comprehensive approaches to fully leverage eye monitoring for cognitive state analysis for many reasons.
@@
The third challenge is the prediction of pupil diameter itself. Previous studies [28, 60, 8] have highlighted that estimating gaze coordinates with a camera involves analyzing images of approximately $60\times 3660 × 36$ pixels [61]. The scale of our images will be smaller for pupil diameter estimation, necessitating analysis at an even finer resolution. This makes accurately predicting pupil diameters more complex than gaze estimation and presents a challenging task.
Figure 1: PupilSense: A web app for estimating and analyzing pupil diameters from everyday images and videos.[A]: Options to select either the left or right pupil for analysis (in blue) and to choose the classification models (in pink). [B]: Visualization of the input and output media, including CAM and estimated pupil diameters. [C]: Estimated pupil diameter values for each frame, analyzed by selected pupil type(s). [D]: EAR values for blink detection, with thresholds for acceptance of open eyes (in green) and rejection (in red). [E]: Consolidated data view showing pupil diameter values, EARs, and differences in pupil diameters, with a downloaded CSV file.
@@
  1. C1
    We provide an open-source webcam-based pupil diameter estimation dataset.
  2. C2
    We propose baseline prediction results using our dataset.
  3. C3
    We implement a novel, user-friendly web application for pupil diameter estimation.
2 Related Work
This section reviews datasets, methods, and applications for gaze and pupil diameter estimation. We highlight how our dataset addresses the gap in publicly accessible pupil diameter resources, offering the largest collection using RGB webcam images and depth maps.
Table 1 compares datasets for gaze and pupil diameter estimation using RGB, RGBD, and IR cameras. While some data were collected in controlled labs, others focused on real-world environments. Most datasets emphasize gaze estimation, highlighting a gap in pupil diameter data, especially in natural settings. Datasets like Rojas et al. [49], Ricciuti et al. [48], and Caya et al. [9] offer valuable contributions to advancing eye-tracking research. However, they are often limited in scope, providing only numerical pupil diameter values instead of images, or are not publicly accessible. In contrast, our dataset is the largest publicly accessible resource for pupil diameter estimation from RGB images taken from webcam images and additionally computed depth maps, contributing significantly to eye-monitoring research.
Table 1: Comparison of related datasets for eye monitoring. While most datasets have gaze coordinates [17, 61, 15, 28, 35, 60, 31, 30, 16, 22, 11], there is a significant gap in pupil diameter informed [49, 48, 9] datasets.
Dataset Subject Size Images Resolution Camera Distance Gaze Vector Public Pupil Diameter
EyeDiap [17] 16 62,500 1920 x 1080 RGBD 80-120 cm 2D, 3D
MPIIFaceGaze [61] 15 213,659 1280 x 720 RGB varying 2D, 3D
RT-GENE [15] 15 122,531 1920 x 1080 RGBD 80-280 cm 3D
Gaze360 [28] 238 172,000 4096 x 3382 RGB varying 3D
SHTechGaze+ [35] 218 165,231 1920 x 1080 RGBD varying 2D
ETH-XGaze [60] 110 1,083,492 6000 x 4000 RGB 100 cm 3D
GW [31] 54 5,800,000 1920 x 1080 IR 0.5-3 cm 2D, 3D
LAEO [30] 485 800,000 variable RGB varying 3D
Fuhl et al. [16] 132 20,867,079 variable IR 0.5-3 cm 2D, 3D
Hou et al. [22] - 35,231 1280 x 720 RGB varying -
Dembinsky et al. [11] 19 648,000 - - 67.5 cm 2D, 3D
Rojas et al. [49] 50 - - - 60 cm 2D, 3D
Caya et al. [9] 16 - variable RGB 10 cm -
Ricciuti et al. [48] 17 20,400 300 x 300 RGB - -
Ours 51 212,073 32 x 16 RGB(D) varying 2D, 3D
Note: The columns of the above table are: (1) the dataset reference (2) the number of subjects; (3) the size of the dataset (4) images available or not, if not, then it implies that only tabular data are available; (5) the resolution of each image; (6) the type of camera(s), our dataset calculates depth after RGB image recordings and hence represents as RGB(D); (7) the distance to the camera(s); (8) type of gaze vector such as 2D or 3D where “D” is a dimention; (9) public dataset or not; and (10) dataset contains pupil diameter or not.
Methods for pupil diameter estimation include PuReST, developed by Santini et al. [50], which tracks pupils robustly using images from 3 head-mounted devices. Ricciuti and Gambi [48] employed video processing with the Viola-Jones algorithm for eye cropping and Canny edge detection with Hough transforms for pupil diameter measurement. Similarly, Caya et al. [9] used preprocessing techniques, including RGB-to-grayscale conversion and the Tiny-YOLO [2] algorithm, achieving percent differences of 0.58% and 0.48% for the left and right eyes, respectively.
Innovative applications for pupil diameter estimation include PupilScreen [39], which uses smartphone cameras in a VR-like enclosure for concussion diagnosis, though its fixed pixel-to-millimeter conversion affects accuracy. Barry et al. [6] employed a smartphone with an external far-red light attachment, while Barry et al. [5] used smartphones’ NIR and RGB cameras with depth calculations for pupil size estimation. Strauch et al. [54] demonstrated pupil diameter as a psychophysiological indicator during video gameplay using an SMI-RED 120 eye-tracker. Studies like Bednarik et al. [7] linked pupillary responses to expertise in microsurgical training, while Palinko et al. [44] and Medathati et al. [40] examined cognitive load and state using pupil diameter in driving simulators and real-world settings. Many of these rely on specialized hardware or close-range setups, limiting accessibility. In contrast, PupilSense offers a device-agnostic, hardware-free platform to democratize pupil monitoring, enhancing accessibility and transparency.
Figure 2: Overview of a data recording and preprocessing (alignment flow). Tobii eye-tracker records pupil diameter, and ChameleonView captures facial recordings using a webcam. Facial recordings start when the participant clicks on the button in the center. The start and end timestamp of the recording is collected in order to synchronize the data with an eye-tracker. To synchronize the 90 frames with the 270 Tobii-captured data points, each metric column is concatenated horizontally across the 90 data points from the three unique timestamps in the Tobii-captured CSV file, followed by computing a row-wise mean.
Figure 3: Pupil diameter distribution of one participant during the recordings. Different pupil diameter measurements and webcam images were captured during the three-second long sessions (in total, 50 sessions). The colors of the boxes indicate the display color used during the recordings (white, black, red, blue, yellow, green, gray, and white again).
3 Data Collection
We created a novel dataset to address the need for high-quality datasets containing eye images with precise pupil diameter annotations in real-world settings. An overview of the data collection process is shown in Figure 2.
We recruited 51 participants (39 males, 11 females, 1 undisclosed, aged 21–44, M=27.58), with consent for the public release of eye-cropped data. Pupil diameter ground truth was recorded at 90 Hz using a Tobii eye-tracker with Tobii Pro Lab software, offering millimeter precision. For video recordings of the face and eyes, we used the built-in webcam of the Microsoft SurfaceStudio 1, which records at a resolution of 1280 × 720 pixels and 30 FPS.
Data was collected using a custom web app, ChameleonView  [^1] https://chameleon-view.netlify.app/, enabling participants to trigger three-second webcam recordings by clicking a central button. As shown in Figure 2, a timestamps file synchronized webcam videos with Tobii data. Each participant completed 50 recordings, with screen background colors varied to capture diverse pupil sizes [57, 47]. The first and last 10 recordings used a white background (#ffffff), while the middle 40 alternated among black (#000000), red (#ff0000), blue (#0000ff), yellow (#ffff00), green (#008000), and gray (#808080). Recordings were conducted in a well-lit room, and the pupil diameter distribution across sessions is visualized in Figure 3.
4 Data Preprocessing
To build the final dataset, we merged the raw webcam footage with Tobii eye-tracker data through two key phases: alignment of recordings and eye cropping.
Figure 4: Data preprocessing pipeline to crop the eyes. For face detection and landmark localization, we used Mediapipe to extract the respective cropped eye images (32x16), left and right, separately. We applied a pre-trained DepthAnythingV2 model on the entire image and cropped the depth maps around the eye regions with the help of landmarks detected from Mediapipe. Next, we applied blink detection on the cropped eyes using the Eye Aspect Ratio (EAR) and a pre-trained vision transformer for blink detection. Cropped eye images and the depth maps are then saved based on the EAR threshold and model confidence score.
4.1 Aligning the Recordings
The process of aligning data from the Tobii eye-tracker and webcam recordings is explained as follows:
@@ 4.1 Aligning the Recordings
    Data sources: Tobii eye-tracker captures pupil diameters and gaze positions at 90 data points per second (90Hz) and webcam recordings capture video at 30 frames per second (30fps).
@@
    Matching timestamps: Each recording has start and end timestamps, see Figure 2 (left). These timestamps are used to extract the corresponding rows from the Tobii eye-tracker data that fall within this time range for synchronization.
@@
    Frame and data count: Each recording is 3 seconds long, resulting in: 90 frames from the webcam (3 seconds $\times$× 30fps) and 270 data points from the Tobii eye-tracker (three seconds $\times$× 90Hz).
  4. 4.
@@
    Aligning frames and data: Concatenate each of the 90 data points from the three unique timestamps and compute a row-wise mean, yielding 90 image frames aligned with 90 data points. And lastly, the first timestamp of the trio is designated as the primary timestamp for that recording, ensuring consistency in data alignment.
This process is repeated for all 50 recording sessions for all 51 participants to ensure uniformity in the dataset.
4.2 Cropping the Eyes
After aligning frames, we crop the eye regions using Mediapipe [38], which detects facial landmarks. To ensure consistency despite variations in participants’ distance from the webcam, eye regions are cropped to fixed dimensions of 32x16 pixels, preserving the natural shape and scale of the eyes.
The full process is presented in Figure 4. Depth maps are generated for the entire image with face, using the DepthAnythingV2 [58] model and cropped based on eye coordinates detected by Mediapipe, allowing us to extract depth information without an RGBD camera. To remove frames with blinks, we use the Eye Aspect Ratio (EAR) calculated from Mediapipe landmarks. Frames with EAR ¡= 0.22 are classified as blinks, while EAR ¿ 0.25 indicates open eyes. A Vision Transformer (ViT) model [12] enhances classification for ambiguous cases (EAR ¿ 0.22 and ¡= 0.25). Frames with blinks are discarded. Given blink durations of 40–200 ms [13, 56], roughly 6 frames per blink are detected at 30fps, making blink removal crucial for data quality.
The final dataset  [^2] https://www.kaggle.com/datasets/vijuls/PupilDiameterDatasets includes 212,073 eye images filtered from 226,912 frames after preprocessing and blink removal. Left and right eye images, along with depth maps, are stored in directories, with a CSV file logging timestamps, session IDs, gaze, pupil data, and frame paths. Sample CSV files, cropped eye images, and depth maps are included in the supplementary materials.
Table 2: Leave one participant out cross validation (LOPOCV) of ResNet18 and ResNet50, evaluated separately for left and right eyes. We excluded one participant per training run and tested the model performance on the left-out participant. This process was repeated for all participants, with the table summarizing the mean and standard deviation of performance metrics across all runs.
Eye Model MAPE $\downarrow$